Please help me find the next number in the sequence,
1,7,31,127,...

Answer by stanbon(75887)

find the next number in the sequence,
1,7,31,127,...
---
It looks like one less than powers of 2.
---
For example:
1st: 1 = 2^(1)-1 = 1
2nd: 7 = 2^(3)-1 = 7
3rd: 31= 2^(5)-1 = 31
4th: 127 = 2^(7)-1 = 127 ...
...
...
nth: 2^(2n-1)-1
-------------------------------
5th: 2^(9)-1 = 511
==========================
Cheers,
Stan H.